# Invariant nature of Raychaudhuri equation in the presence of Inflaton fields

### v1

@@ Abstract
The aforementioned solution corresponds to $H^{2}\simeq\frac{1}{9t^{2}}$, which directly follows $a\propto t^{1/3}$ and $\rho\propto a^{-6}$. This solution is exact for a massless scalar field. At this stage, the field’s trajectory rapidly ascends and eventually reaches the attractor. This greatly expands the range of initial conditions that result in an inflationary phase. Once the trajectory intersects the attractor at a point when it is flat, denoted as $|\phi|>>1$, the subsequent solution elucidates a phase of accelerated expansion. Then the corresponding scalar field can be written as
@@
where $t_{i}$ is the time when the expansion starts and $t_{f}$ is the time when the scalar field dies out. It should be noted that the solution obtained in (20) in our particular setting exhibits a comparable linearly time-dependent solution of $\varphi$. Hence, it may be posited that the K-essence scalar field, which maintains the invariance of the Raychaudhuri equation, is similar to an “inflaton” field.
Case-II: Inhomogeneous field: Arguments for the inflationary scenario are well given in Mukhanov [43]. We just state that the perturbations that occur during inflation are the source of the inhomogeneities; for this reason, these perturbative fields ($\Phi$) and ($\delta\varphi$) are referred to as inflaton fields. In Ref. [43], they have used the same type of action (1), the gauge-invariant metric perturbation ($\Phi$) and scalar field perturbation ($\delta\varphi$) has been expressed as
@@
$\delta\varphi$ $\cong C\sqrt{\frac{1}{c_{s}p_{X}}}(\pm ic_{s}\frac{k}{a}+H+...)\leavevmode\nobreak\ exp(\pm ik\int\frac{c_{s}}{a})dt\leavevmode\nobreak$ (28)
which oscillate in the short-wavelength regime. Here, $C$ is an integration constant, $p(X)\equiv\mathcal{L}(X)$, $p_{X}\equiv\frac{\partial p}{\partial X}=\mathcal{L}_{X}$ and $\dot{\varphi_{0}}$ is the square root of the kinetic part of the homogeneous scalar field.
Now we consider DBI type Lagrangian as $p(X)=\mathcal{L}(X)=1-V\sqrt{1-2X}$ and $B(X)=4\pi C\dot{\varphi}_{0}\sqrt{\frac{p_{X}}{c_{s}}}$. Then
$B(X)\cong D\Big{(}1+X+\frac{3}{2}X^{2}\Big{)}$ (29)
[…]
$\exp{(iA(t))}=1+iA(t)-\frac{A^{2}(t)}{2!}+\frac{(iA(t))^{3}}{3!}-...$ (30)
Therefore, for real solutions, we can recast the above inhomogeneous inflaton fields (27) and (28) as
$\Phi\simeq D\Big{(}1+X+\frac{3}{2}X^{2}+\frac{5}{2}X^{3}\Big{)}\Big{(}1-\frac{A^{2}(t)}{2!}-\frac{A^{4}(t)}{4!}\Big{)}$ (31)
$\delta\varphi$ $\simeq C\sqrt{V}\Big{[}-\frac{k}{a}\Big{(}A(t)-XA(t)-\frac{X^{2}}{2}A(t)+\frac{A^{3}(t)}{3!}$ (32)
$-X\frac{A^{3}(t)}{3!}-\frac{X^{2}A^{3}(t)}{3!}\Big{)}+H\Big{(}1-\frac{A^{2}(t)}{2!}-\frac{A^{4}(t)}{4!}\Big{)}\Big{]}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$
@@
where $\mathcal{D}=-\frac{w}{H_{0}^{3/2}}$ and $\mathcal{A}(t)=\frac{H_{0}a^{2}}{w^{2}}$. Thus far, we have expressed our general solution of field (23) in terms of $X$ and $\mathcal{A}(t)$, while ensuring that the constraint $v^{\mu}\partial_{\mu}X=0$ is maintained to ensure the invariance of the Raychaudhuri equation (11) without considering the curvature effect, namely the Ricci tensor of gravity. The solution of the field $\phi$ (34) carries some resemblance to the inhomogeneous inflation fields (31) and (32) under certain conditions for the DBI type action mentioned earlier. It should be noted that the matching of our solution (34) with the inflaton fields Eq.s (31) and (32) requires a discussion of the conditions. However, we cannot explore this topic here, as the inhomogeneous solutions of the fields have different origins and depend on the specific model Lagrangian $\mathcal{L}(X)$ or $p(X)$ that we have employed. Hence, the inhomogeneous fields we have described in Eq.s (23) or (34) can be referred to as inflaton fields, even though they do not originate from any cosmological analysis in the early universe. Therefore, the Raychaudhuri equation can be employed to elucidate the origins of the primordial inhomogeneities via fields associated with inflation.
Concluding remarks: In the presence of both homogeneous and inhomogeneous scalar fields, the modified Raychaudhuri equation (11) returns to its original form. The Raychaudhuri equation is invariant in this sense. This may be conceivable since the Raychaudhuri equation is based on geometry and not on any specific theory of gravitation. Based on our findings, the universe is likely inhomogeneous. Remember that the Raychaudhuri equation has a geometrical identity and should be dynamical if we incorporate the Ricci tensor component from any gravity theory. Thus, the modified Raychaudhuri equation in its standard form is invariant for all gravitational theories, regardless of scalar field homogeneity or inhomogeneity. The Guardian [44] and recent observations [45, 46, 47] indicate that the cosmos is inhomogeneous. This observational evidence is in sync with our results obtained from the equation regarding inhomogeneous scalar fields. It is vital to highlight that this research does not analyze the universe’s cosmic dynamics. Such investigations of cosmic dynamics for homogeneous scalar fields in the light of the Raychaudhuri equation has been done in [12].
On the other hand, we consider the scale factor (17) for the solution of Eq. (16), which may be used for the late-time acceleration era or inflationary era of the universe [41, 42, 40, 43]. The inflationary paradigm predicts large-scale spatial homogeneity and isotropy in the cosmos. This suggests that the cosmos is the same everywhere. Other scenarios employing an inhomogeneous scalar field driving inflation with spatially different properties are conceivable. Inhomogeneities in the early universe may have been caused by quantum fluctuations and other mechanisms. These inhomogeneities may affect inflation predictions and density fluctuations. Inhomogeneous inflaton fields are caused by quantum perturbations, which is the main source of primordial inhomogeneities [43]. According to Mukhanov [43], we can say that our case’s homogeneous and inhomogeneous fields also resemble inflaton fields under specific circumstances. Thus, the Raychaudhuri equation may be utilized to study primordial inhomogeneities utilizing inflation-related factors.
As a final comment, in this work, we solely focused on the invariance property of the Raychaudhuri equation in the presence of scalar fields, which can be called inflaton fields. We did not consider the cosmological behaviors and other effects of the universe.

### v2

@@ V Inflaton fields:
This corresponds to $H^{2}\simeq\frac{1}{9t^{2}}$, which directly follows from $a\propto t^{1/3}$ and $\rho\propto a^{-6}$. This solution is exact for a massless scalar field. The field’s trajectory rapidly ascends and eventually reaches the attractor. The range of initial conditions increase and result in an inflationary phase. Once the trajectory intersects the attractor at a point when it is flat, denoted as $|\varphi|>>1$, the subsequent solution elucidates a phase of accelerated expansion. Then the corresponding scalar field is
@@
where $t_{i}$ is the time when the expansion starts and $t_{f}$ is the time when the scalar field dies out. Our solution (19) in our particular setting exhibits a comparable linearly time-dependent solution for $\phi$. Hence, the K-essence scalar field that maintains the Raychaudhuri equation’s invariance is similar to an “inflaton” field.
Case-II: Inhomogeneous field:
Next, we show that our inhomogeneous solution (22) is similar to the inflaton solutions Eq.s (8.65) and (8.66) of Mukhanov [38] under specific circumstances using DBI type Lagrangian.
@@
$\Phi\simeq D\Big{(}1+X+\frac{3}{2}X^{2}+\frac{5}{2}X^{3}\Big{)}\Big{(}1-\frac{A^{2}(t)}{2!}-\frac{A^{4}(t)}{4!}\Big{)}$
$\delta\varphi$ $\simeq C\sqrt{V}\Big{[}-\frac{k}{a}\Big{(}A(t)-XA(t)-\frac{X^{2}}{2}A(t)+\frac{A^{3}(t)}{3!}$
$-X\frac{A^{3}(t)}{3!}-\frac{X^{2}A^{3}(t)}{3!}\Big{)}+H\Big{(}1-\frac{A^{2}(t)}{2!}-\frac{A^{4}(t)}{4!}\Big{)}\Big{]}$
@@
where $\mathcal{D}=-\frac{w}{H_{0}^{3/2}}$ and $\mathcal{A}(t)=\frac{H_{0}a^{2}}{w^{2}}$. So far, we have expressed our general solution of field (22) in terms of $X$ and $\mathcal{A}(t)$, while ensuring that $v^{\mu}\partial_{\mu}X=0$, maintaining the form invariance of the Raychaudhuri equation (10) without considering the curvature effect, namely the Ricci tensor of gravity. The solution of the field $\phi$ (29) carries some resemblance to the inhomogeneous inflation fields (26) and (27) under certain conditions for the DBI type action mentioned earlier.
Thus, the Raychaudhuri equation becomes relevant when considering primordial inhomogeneities within the framework of inflationary cosmology. In the inflationary scenario, the universe experiences a rapid exponential expansion, effectively smoothing out any inhomogeneities and anisotropies.
However, during inflation, small inhomogeneities [38] can still remain that eventually develop into the large-scale structures we observe in the present-day universe [47, 48, 49, 50]. So that the Raychaudhuri equation may be employed to elucidate the origins of the primordial inhomogeneities via fields associated with inflation.
Alternatively, our particular inhomogeneous solutions (21), (22), and (28) of the K-essence scalar fields derived from the invariant Raychaudhuri equation (10) can be relevant when studying inhomogeneous cosmology which can be mapped on to the Lemaitre-Tolman-Bondi (LTB) metric [53, 54, 55, 56, 57]. In a different context, Gangopadhyay et al. [20] developed the scaling relation for inhomogeneous K-essence scalar fields $\phi(r,t)=\phi_{1}(r)+\phi_{2}(t)$ (14) in the context of the LTB metric.
VI Concluding remarks:
We have shown that the modified Raychaudhuri equation (10) reverts to its original form in the presence of both homogeneous and inhomogeneous scalar fields which are similar to inflatons. This is conceivable because the Raychaudhuri equation is based on geometry and not on any specific theory of gravitation. Our solutions based on the Raychaudhuri equation admit an inhomogeneous universe as well as a homogeneous universe. So the Raychaudhuri equation is compatible with both homogeneous and inhomogeneous universe. We recall that there exists another interesting form invariance in cosmology, viz., form invariance of the metric tensor, which leads to the existence of killing vectors.
Sarkar [47] has provided a comprehensive explanation of inhomogeneous cosmology [48], supported by empirical data. In this article, the author elucidated the need of a novel kind of structure development for inhomogeneous and anisotrpic universes, as discussed by Krasinski [48]. However, these models have not been extensively investigated, particularly in terms of their implications for structure formation, in comparison to the FLRW models. Recent observations [49, 50, 51, 52] indicate that the cosmos is inhomogeneous. This observational evidence is in sync with our results obtained from the equation regarding inhomogeneous scalar fields. Investigations for homogeneous scalar fields in light of the Raychaudhuri equation have been done in [7].
@@ VI Concluding remarks:
Conflicts of interest: The authors declare no conflicts of interest.
Data availability: There is no associated data with this article, and as such, no new data was generated or analyzed in support of this research.